OpenStudy (anonymous):

Which of the following expressions are equivalent? Justify your reasoning. OPTIONS LISTED BELOW

OpenStudy (anonymous):

A. \[\sqrt[4]x{3}\] B. \[\frac{ 1 }{ x^{-1}}\] C.\[\sqrt[10]{x}^5 \times x^4 \times x^2\] D. \[x^{1/3} \times x ^{1/3} \times x^{1/3}\]

OpenStudy (triciaal):

review exponent rules see which applies solve and choice what matches

OpenStudy (triciaal):